A line L is perpendicular to the line 5x-y=1 and the area of triangle formed by the line L and coordinate axes is 5.The equation of line L is?

Grade:11

1 Answers

Rhythm Jain
6 years ago
Hi, According to question L is perpendicular to the line 5x-y=1. Now the equation of L becomes y+5x=d (here d is a varaiable) which gives y/d+5x/d=1. Now using intercept form of straight line we get the x=d/5 and y=d so the coordinates of the L becomes (0, d) and (d/5, 0). Now area of triangle formed is 1/2*d*d/5=5 => d^2/10=5 => d=sqrt(50)